What is the probability that sample mean is more than 12 million
1
Expert's answer
2020-12-13T17:59:00-0500

P("\bar x >12)"

Assuming that the sample size, population mean and standard deviation are known,

Z="\frac {\bar x - \mu} {\frac{ \sigma} {\sqrt n}}"

= "\frac {12 - \mu} {\frac{ \sigma} {\sqrt n}}"

Read the value of "P(z>\frac {12 - \mu} {\frac{ \sigma} {\sqrt n}} )" from z-table.